# DEER: A Comprehensive and Reliable Benchmark for Deep-Research Expert Reports

### v1

@@ 4.4 Information Verification
To automate fact verification in expert reports, we must overcome the limitations of existing systems that focus solely on explicit citations. In reality, many sentences implicitly inherit preceding evidence or construct new claims through contextual connections. Excluding these implicit claims from verification makes the reliability assessment of the entire document incomplete. Our Information Verification Module proposes an architecture that extends the verification scope to the entire document by tracking sentence-level logical flows.
@@
When processing long documents as a single input, LLMs suffer from the Lost-in-the-Middle problem, which leads to the omission of important claims liu2023lostmiddlelanguagemodels. To address this, we apply a Batch Extraction strategy that maintains the context of the entire document $D$ while parallel-processing only the target sentences divided into small batches $B_{j}$:
@@
The extracted claims are passed to the fact verification stage, where the biggest technical challenge is the trade-off between accuracy and cost. Verifying every claim individually yields high accuracy, but cost increases linearly with long documents. To address this, we introduce two scaling strategies.
First, we apply Grouped Claim Verification, which bundles multiple claims citing the same source into a single input for verification. Experimental results showed that increasing Group Size from 1 to 10 raised accuracy from 0.78 to 0.89 and reduced cost per claim by 85% (Section 6.5). This indicates that grouped input helps in understanding the overall structure of the document and verifying claims, while reducing verification overhead. Second, greater API cost reduction is possible through the Context Retrieval strategy, which selectively provides only the context necessary for fact verification. This method achieved an API cost reduction of approximately 24 times, but a 6% drop in accuracy was observed, indicating a clear trade-off.
This hierarchical pipeline (Batch Extraction $\rightarrow$ Grouped Verification $\rightarrow$ Context Retrieval) is designed to achieve both high-precision verification and cost efficiency even in long reports containing hundreds of claims.
@@ 5 Experiment Setup
5.1 Implementation Details.
We use the LLM-as-a-judge method to evaluate report quality. The evaluation model takes as input the task query, the report to be evaluated, expert guidance, and detailed rubrics and evaluation criteria. The evaluation is conducted across 5 dimensions: Request Completeness, Evidence Validity, Structure Consistency, Narration Style, and Ethics Compliance, with each rubric item scored on a 1–10 scale. The model generates a score and a short rationale justifying that score for each item, and the final output is in JSON format consisting of (rationale, score) pairs for each item.
Information Integrity and Information Sufficiency are evaluated through a separate verification pipeline, the Information Verification Module. This module extracts claims from the report and assesses their factuality by comparing them with external evidence obtained via web searches. The verification results are calculated as detailed metrics such as Claim Factuality and Citation Validity.
The evaluation is hierarchical, computing the final quality score by sequentially aggregating the averages of lower-level rubric items to higher levels, using the scores collected above.
Model Req. Comp. Evid. Valid. Struct. Cons. Narr. Style Info. Int. Info. Suff. Ethics Mean
General LLM
qwen3-235b-fast 6.10 6.59 9.04 8.66 3.30 4.40 8.58 6.67
gemini-2.5-pro-fast 6.08 7.02 9.29 8.87 2.92 4.43 8.89 6.79
claude_opus4.1_fast 6.04 6.47 8.96 8.95 4.76 4.81 8.78 6.97
gpt5_fast 5.61 6.63 8.80 8.46 2.18 3.45 8.82 6.28
LLM+Thinking
qwen3-235b-think 6.77 6.53 9.32 8.92 3.28 4.39 8.75 6.85
gemini-2.5-pro-think 6.41 7.66 9.55 9.20 3.86 4.97 9.37 7.20
claude_opus4.1_think 5.93 6.41 9.05 9.04 4.72 5.12 8.99 7.04
gpt5_think 7.69 8.37 9.57 9.32 4.19 5.42 9.37 7.70
LLM+Thinking+Search
qwen3-235b-think_search 6.43 6.31 9.07 8.82 3.60 4.49 8.44 6.74
gemini-2.5-pro-think_search 6.61 7.58 9.50 9.11 4.62 5.02 8.97 7.34
claude_opus4.1_think_search 6.03 6.62 9.04 9.03 5.17 4.80 8.85 7.08
gpt5_think_search 7.88 8.26 9.39 9.04 6.42 6.44 9.38 8.12
Deep Research
webthinker 5.27 5.76 8.48 8.62 7.57 6.92 8.68 7.33
qwen3-235b-deep 4.95 6.00 7.28 8.31 7.13 8.29 8.66 7.23
gemini-2.5-pro_deep 6.08 6.81 8.81 8.78 7.13 7.87 8.62 7.73
claude_opus4.1_deep 6.60 6.91 8.90 8.85 6.79 7.66 8.75 7.78
gpt5_deep 6.26 7.11 9.20 8.93 7.64 7.13 8.99 7.89
Table 2: Baseline models’ expert report evaluation results. In each column, the highest score is marked in bold and the second highest score is underlined.
@@ 5.2 Baseline Models
In this study, to compare the expert report generation capability of LLM-based systems, we established four categories of baseline models. Each category is designed to increase in complexity incrementally, ranging from simple language model response generation to specialized research systems that combine search and deep reasoning.
(1) General LLMs This is the basic LLM setting that generates responses in a single pass (inference), serving as a baseline for measuring response quality in language-generation-based models. This category includes qwen3-235b-fast qwen3technicalreport, gemini-2.5-pro-fast google2025gemini, claude_opus4.1_fast anthropic2025claude, and gpt5_fast openai2025gpt5system models.
(2) LLMs with Reasoning This setting uses the reasoning function on the same LLMs as in (1), extending the internal reasoning step before answer generation. This category includes qwen3-235b-think, gemini-2.5-pro-think, claude_opus4.1_think, and gpt5_think models.
(3) LLMs with Reasoning + Web Search This configuration combines the LLMs with the Reasoning setting and web search, evaluating scenarios in which external information is used during reasoning. This category includes qwen3-235b-think_search, gemini-2.5-pro-think_search, claude_opus4.1_think_search, and gpt5_think_search models.
(4) Deep Research Frameworks (Open-Source & Commercial) This is the upper baseline, comprising specialized deep research frameworks and commercial research support systems. We used webthinker webthinker as the open-source framework, and commercial systems include qwen3-235b-deep, gemini-2.5-pro_deep, claude_opus4.1_deep, and gpt5_deep. Detailed configurations and hyperparameters for all baseline models are summarized in Appendix A.
@@ 6.1 Main Results
Summarizing Table 2, all models show generally high performance in metrics related to form, style, and ethics, such as Structure Consistency (Struct. Cons.), Narration Style (Narr. Style), and Ethics Compliance (Ethics). However, they consistently show relatively lower scores on metrics that assess how faithfully they address the requirements in the user’s query and provide persuasive evidence in terms of calculation, methodology, and logical development, such as Request Completeness (Req. Comp.) and Evidence Validity (Evid. Valid.). This indicates that while the latest LLMs are already quite proficient in structuring reports, writing naturally, and adhering to basic ethical standards, their ability to accurately meet user intent at an expert level and support it with sufficiently robust argumentation and evidence structures remains limited. Reasoning models showed consistent improvements across almost all evaluation metrics relative to General LLMs, suggesting that simply expanding the reasoning budget without additional tools improves coverage of requirements, logical development, evidence selection, and narrative consistency. Models combining this with web search show additional gains in information-related metrics. The Deep Research category consistently achieves higher scores in Information Integrity (Info. Int.) and Information Sufficiency (Info. Suff.) than other categories, providing the strongest baseline for the broad collection and organization of external documents.
However, it is characteristic that the Deep Research category consistently records lower scores than reasoning-based single LLMs in the remaining key evaluation axes, especially Req. Comp. and Evid. Valid. These results suggest that the large amount of external material injected through search in the Deep Research configuration may cause the reference point for report development to gradually shift from the query itself to external knowledge, inducing a kind of retrieval-induced drift in which information rather than blurs the problem definition and argumentation structure. Therefore, this demonstrates that simply increasing the volume of search and citations does not guarantee the quality of expert reports.
6.2 Fine-grained Analysis of Evaluation Results
Figure 3(a) decomposes the upper evaluation dimensions into detailed items to summarize the performance patterns of current Deep Research and LLM-based systems in expert reports. In Request Completeness, scores across sub-items are generally low, with scope the lowest, suggesting a lack of clarity about what the report covers and excludes, and under what assumptions and limitations the discussion is developed. In Evidence Validity, the score for the numeric item is slightly lower than that for logical, indicating that current systems make mistakes relatively more frequently in terms of quantitative accuracy, such as calculation processes, used formulas/statistical models, and selection of indicators/units, rather than logical development itself. In Information-related detailed items, ref_diversity and refs_amount appear commonly low, meaning that even if the overall structure and narrative are compliant, the number and diversity of sources actually referenced are limited.
Figure 3(b) shows the domain-specific performance of Deep Research agents as the average of two randomly selected tasks in each domain (domain-specific average scores for all tasks are presented in Appendix LABEL:app:domain-results). While most domains are distributed in a relatively similar range around 7 points, in the case of Physics, most Deep Research agents except gpt5_deep consistently record the lowest scores compared to other domains, and relatively large deviations between Deep Research agents also appear in CS, History, and Linguistics.
@@ 6.3 Correlation with Human Evaluation
To verify the validity of the proposed evaluation method and analyze the contribution of each component, we conducted experiments to assess its correlation with human expert evaluation. For 15 tasks across 5 domains, 3 reports randomly selected from those generated by the Deep Research systems used in 6.1 were used as evaluation targets. Each report was evaluated by two independent experts, collecting a total of 90 human evaluation samples. To assess the alignment between the evaluation method and human experts, we used the Pearson correlation coefficient, the Spearman rank correlation coefficient, and pairwise agreement deepbench. Pairwise agreement is a metric that measures the proportion of cases in which our evaluator chose the same priority as the human expert when comparing two reports, indicating how consistently it reproduces human relative judgments across tasks. Evaluations were performed on the same reports using 5 LLM-based evaluation models (GPT-5, GPT-5-mini, Claude Opus 4.1, Claude Sonnet 4.5, Gemini 2.0 Pro), and the correlation between the average of the 5 model evaluations and the human evaluation was analyzed.
@@
Vanilla 0.64 0.61 0.66
+ Dimensions 0.67 0.65 0.80
+ Granular Checklist 0.53 0.46 0.72
+ Expert Guidance 0.75 0.71 0.84
@@
Table 3: Average human correlation across five AI models with incremental addition of evaluation components. Inter-Human shows inter-annotator agreement.
Table 3 shows how the correlation with human evaluation changes as evaluation components are added step by step. Starting from the Vanilla deepbench setting, which judges overall quality without explicit evaluation criteria, we sequentially expanded by introducing evaluation dimensions (+Dimensions), decomposing them into granular rubrics (+Granular Rubric), and finally adding task-specific expert guidance (+Expert Guidance) to measure the impact of each component. The Vanilla setting showed significant correlation with human experts even without explicit criteria, but this correlation decreased as evaluation dimensions (+Dimensions) were introduced and decomposed into detailed items (+Granular Rubric). Since report evaluation is a high-dimensional task that requires simultaneous consideration of various elements, such as logic, concepts, analysis, and domain knowledge, across long and complex texts, each rubric item comprises broad, interpretive judgment units rather than narrow rules. It is analyzed that the increase in the number of items led to an increase in interpretive variability rather than refinement, resulting in a drop in human correlation. Adding task-specific expert guidance (+Expert Guidance) clarifies the interpretation method for each criterion, restoring consistency and significantly improving correlation across all metrics. Writing longer rationales (+Longer Rationale) further improves correlation by enriching the basis for each item. In conclusion, DEER tends to achieve relatively stable criterion application and high alignment with human experts, even in report evaluation settings that must consider multiple evaluation dimensions simultaneously.
6.4 Inter-evaluator Agreement (IEA)
To verify that the proposed evaluation method yields consistent results across different LLMs, we measured inter-evaluator reliability among the 5 evaluation models used previously IAA; checkeval. We measured the agreement between models using Krippendorff’s $\alpha$ and ICC (Intraclass Correlation Coefficient). High agreement suggests that the proposed method possesses versatility and stability that are not dependent on a specific LLM.
@@ 6.4 Inter-evaluator Agreement (IEA)
Table 4 shows how the agreement between evaluation models changes as evaluation components are added step by step. The Vanilla setting showed an appropriate level of inter-model agreement, but agreement decreased significantly when introducing evaluation dimensions (+Dimensions) and decomposing them into detailed items (+Granular Rubric). As explained earlier, this is because rubric items consist of interpretive judgment units, so the increase in the number of items increased the variability in interpretation across models. Adding expert guidance (+Expert Guidance) significantly improved agreement across all metrics, restoring evaluation consistency. In particular, the variation focusing only on the top rubric items achieved the highest agreement. This shows that expert guidance clarifies the interpretation of evaluation criteria, enabling consistent evaluation even between different LLM evaluation models.
@@ 6.5 Information Verification Module Evaluation
In this section, we evaluate the proposed Information Verification Module by dividing it into two stages: claim extraction and fact verification.  [^2] Detailed content is covered in Appendix F.
@@
We compared and analyzed 728 claims identified by human experts (Ground Truth) and LLM extraction results using a Paragraph-Level LLM Judge (Table 5). As a result of measuring semantic Recall beyond simple count comparison, GPT-5-mini showed significantly higher performance compared to GPT-4.1. In particular, the GPT-5-mini (Batch 5) setting recorded a Recall of 0.9333 for Type A, B, and C claims requiring external evidence, capturing human expert claims almost perfectly. Also, the number of claims extracted per paragraph (Claims/Para) was the highest at 6.95, confirming that detailed information in the document is extracted densely without omission. On the other hand, GPT-4.1 showed relatively lower Recall (about 0.83) and density (about 4.8), suggesting a tendency for the model to excessively summarize information or omit some details. In conclusion, for precise fact-checking, a strategy of maximizing recall by operating high-performance models like GPT-5-mini with small batch sizes (Batch Size 5, 10) is essential.
Table 5: Comparison of Claim Extraction Performance across Models and Batch Sizes. Recall is measured at the Paragraph-Level using LLM Judge.
Model Batch Size Recall Claims/Para
Ground Truth - - 7.22
GPT-4.1 5 0.9057 5.00
10 0.8897 4.84
20 0.8897 4.65
GPT-5-mini 5 0.9333 6.95
10 0.9264 6.22
20 0.9218 6.05
@@
In the claim verification stage, we conducted an ablation study on grouping and retrieval settings using GPT-4.1 (Table 6). When applying Grouped Claim Verification without retrieval, the F1 score remained comparable (91 → 90), while the cost per claim decreased by approximately 82%, demonstrating that batching claims provides substantial efficiency gains with minimal performance loss.
Additionally, enabling retrieval further reduced the cost to \$0.004 per claim but lowered the F1 score to 85. Therefore, for scenarios requiring high reliability, grouped verification without a retrieval configuration offers the best balance, while with a retrieval setting, it is preferable when maximizing cost efficiency is the primary objective.
Grouped Retrieval F1 Cost (\$/Claim)
0.91 0.083
20 0.90 0.015
20 0.85 0.004
Table 6: Ablation study on GPT-4.1. Grouping significantly improves cost efficiency with minimal F1 degradation; retrieval further reduces cost at the expense of accuracy.
@@ 7 Conclusion
We introduce DEER, a comprehensive benchmark for evaluating Deep Research agents, featuring a multi-dimensional taxonomy and a hybrid evaluation framework that combines LLM-based qualitative assessment with a retrieval-augmented Information Verification Module. Our experiments reveal that while state-of-the-art models excel in structure and style, they still lag behind human experts in critical areas like evidence validity and information sufficiency, although "Deep Research" systems show promising improvements. By demonstrating a high correlation with human judgments, DEER offers a scalable standard for advancing autonomous research agents toward trustworthy, expert-level performance.
@@
Acknowledgments
Benchmark Human rubrics Expert curated Open- ended Non-tech domains LLM judge Claim fact-check Inter- pretability Systematic criteria Avg. rubrics
AcademicBrowse AcademicBrowse
BrowseComp browsecomp
Mind2Web2 mind2web2 $\triangle$ 50
ResearchBench ResearchBench
ResearcherBench ResearcherBench $\triangle$ $\triangle$ 14
LiveDRBench characterizing $\triangle$
SPOT SPOT $\triangle$
ExpertLongBench expertlongbench 16
ReportBench ReportBench $\triangle$
DeepScholar-Bench DeepScholar-Bench $\triangle$ $\triangle$
DeepResearchBench deepbench $\triangle$ $\triangle$ 25
LiveResearchBench LiveResearchBench $\triangle$ $\triangle$
DeepResearchArena DeepResearchArena $\triangle$ $\triangle$
DeepResearchGym deepgym $\triangle$
ResearchRubrics ResearchRubrics $\triangle$ 26
DREX (Ours) 130+
Table 7: Comparison of DREX with representative deep research benchmarks. Here, indicates full support, no support, and $\triangle$ partial support. For Claim fact-check, $\triangle$ means that only a subset of claims (e.g., explicitly cited or gold-labeled ones) are checked rather than all explicit and implicit claims. For Interpretability, $\triangle$ means that the benchmark provides only coarse, dimension-level insight (e.g., a few high-level scores) without shared checklist-level diagnostics across tasks.

### v2

@@ 4.4 Information Verification
To automate information verification in expert reports, we must overcome the limitations of existing systems that focus solely on explicit citations. In reality, many sentences implicitly inherit preceding evidence or construct new claims through contextual connections. Excluding these implicit claims from verification makes the reliability assessment of the entire document incomplete. Our Information Verification Module proposes an architecture that extends the verification scope to the entire document by tracking sentence-level logical flows.
@@
When processing long documents as a single input, LLMs suffer from the Lost-in-the-Middle Liu et al. (2023a) problem, which leads to the omission of important claims. To address this, we apply a Batch Extraction strategy that maintains the context of the entire document $D$ while parallel-processing only the target sentences divided into small batches $B_{j}$:
@@
The extracted claims are passed to the fact verification stage, where the biggest technical challenge is the trade-off between accuracy and cost. Verifying every claim individually yields high accuracy, but cost increases linearly with long documents. To address this, we introduce two scaling strategies. First, we apply Grouped Claim Verification, which bundles multiple claims citing the same source into a single input for verification. Experimental results showed that increasing Group Size from 1 to 20 raised F1 from 87.25 to 91.61 and reduced cost per 1K claims by 72%. This indicates that grouped input improves understanding the overall structure of the document and verifying claims, while reducing verification overhead. Second, greater API cost reduction is possible through the Context Retrieval strategy, which selectively provides only the context necessary for fact verification. This method achieved an API cost reduction of approximately 72%, but a 3.4% drop in F1 was observed, indicating a clear trade-off. This hierarchical pipeline (Batch Extraction $\rightarrow$ Grouped Verification $\rightarrow$ Context Retrieval) is designed to achieve both high-precision verification and cost efficiency even in long reports containing hundreds of claims.
@@ 5 Experiment Setup
5.1 Implementation Details
We evaluate report quality using an LLM-as-a-judge approach across five dimensions—Request Fulfillment, Analytical Soundness, Structural Coherence, Format & Style, and Ethics Compliance. The LLM judge receives the task query, the report being evaluated, task-specific Expert Evaluation Guidance, and a fixed set of fine-grained rubric items. It assigns a 1–10 score and a brief rationale to each item and returns the results as a JSON object mapping each item to its score and rationale. In parallel, Information Integrity and Information Sufficiency are assessed by the Information Verification Module, which extracts claims from the report and verifies them against external evidence retrieved via web search, producing metrics such as Claim Factuality and Citation Support. Details are provided in Appendix G.4. Finally, we compute the overall quality score hierarchically by aggregating scores bottom-up from rubric items to higher-level categories. Details of the scoring procedure are provided in Appendix E.3.
@@ 5.2 Baseline Models
To compare expert report generation performance, we consider four baseline families: General LLMs (fast), LLMs+Reasoning (think), LLMs+Reasoning+WebSearch (think+search), and Deep Research (deep). Each family is instantiated using multiple model families/backbones (Qwen, Gemini, Claude, and GPT), and they differ in whether they (i) use reasoning, (ii) use web search, and (iii) employ research-system orchestration. Detailed model configurations are provided in Appendix H.
@@ 6.1 Main Results
Summarizing Table 1, most models score highly on structure, style, and ethics (Structural Coherence, Format & Style, Ethics), but lag on requirement fulfillment and analytical soundness (Req. Fulfillment, Analytical Soundness). This suggests that report-writing quality is largely mature, whereas expert-level intent alignment and evidence-based justification remain limited. Across baseline families, adding reasoning in think improves overall performance over fast. Moreover, external-information metrics—Information Integrity and Information Sufficiency—see further gains with think+search and deep. An interesting finding is that reasoning models without web search (think) outperform think+search and deep on report-quality metrics excluding information-related scores. This suggests that integrating diverse external information can blur the problem definition and argument structure. Accordingly, simply adding search and external sources may not guarantee improved report-writing quality.
6.2 Fine-grained Analysis
Figure 3(a) breaks down the performance patterns of Deep Research systems on expert report generation into fine-grained dimensions. The sub-dimension scores for Request Fulfillment, Analytical Soundness, and Structural Coherence are generally low, with scope particularly low within Request Fulfillment. This indicates that Deep Research systems often fail to clearly specify the report scope (what is covered vs. excluded) and the assumptions/constraints used to develop the report. In addition, ref_amount is low in Information Sufficiency, suggesting that systems tend to rely on a small number of references rather than leveraging a broad set of sources. Figure 3(b) shows domain-level performance of Deep Research agents. For each domain, we report results on two tasks randomly selected from that domain. Performance tends to drop uniformly across agents on relatively difficult tasks, and Physics consistently records the lowest scores compared to other domains.
@@ 6.3 Correlation with Human Evaluation
Vanilla 0.64(0.16) 0.61(0.17) 0.66(0.14)
+ Dimensions 0.67(0.10) 0.65(0.14) 0.80(0.07)
+ Granular Rubrics 0.62(0.14) 0.59(0.17) 0.78(0.08)
+ Expert Guidance 0.75(0.07) 0.71(0.06) 0.84(0.03)
@@
Table 2: Average human correlation across five AI models with incremental addition of evaluation components (reported as mean(std)). Inter-Human shows inter-annotator agreement.
To validate the proposed evaluation method and assess the contribution of each component, we compared LLM-based evaluations against human expert judgments. For each of the 45 reports, we collected two independent ratings from domain experts whose expertise aligned with the report’s topic (90 ratings total). We computed Pearson’s r, Spearman’s ρ, and LLM–human pairwise agreement for each of the five LLM-based evaluator models, and report the average of each metric across models. Additional details on the experimental setup and human evaluation protocol are provided in Appendix I.
Table 2 shows how alignment with human judgments changes as evaluation components are added step by step. Vanilla, which assesses overall quality holistically, and +Dimensions, which introduces high-level evaluation dimensions, both achieve relatively high correlation with human judgments. However, Pearson’s r and Spearman’s ρ drop at the +Granular Rubrics stage where the rubric is further decomposed into many fine-grained items. In contrast, task-specific +Expert Guidance helps evaluators apply these rubric items by surfacing domain-relevant cues that non-experts may overlook, thereby achieving the highest correlation with human judgments.
6.4 Inter-evaluator Reliability
@@
To verify that the proposed evaluation method yields consistent results across different LLM evaluators, we measure inter-evaluator reliability Artstein (2017); Lee et al. (2025). Specifically, we compute Krippendorff’s $\alpha$ and the intraclass correlation coefficient (ICC) using the scores from the five LLM judges described in Section 6.3.As shown in Table 3, Vanilla achieves moderate inter-evaluator reliability, but reliability drops substantially under +Dimensions. +Granular Rubrics yields a slight recovery, though still below Vanilla. +Expert Guidance attains the highest reliability across metrics. This suggests that expert guidance clarifies what to look for under each criterion, enabling more consistent judgments across different LLM judges.
@@ 6.5 Information Verification Module Evaluation
In this section, we evaluate the proposed Information Verification module by dividing it into two stages: claim extraction and claim verification.  [^2] Detailed content is covered in Appendix F.
@@
Table 4 compares claim extraction performance across models and batch sizes using paragraph-level semantic recall measured by an LLM Judge. GPT-5-mini achieves the highest recall (92.17%) and classification F1 (68.80) at batch size 10, while incurring substantially lower cost than GPT-5. Although GPT-5 extracts more claims per paragraph, its recall and F1 saturate, indicating that denser extraction does not improve coverage. Increasing the batch size consistently reduces cost while incurring only moderate degradation in recall, making GPT-5-mini with a batch size of 20 a cost-effective configuration for large-scale claim extraction.
Model Batch Density Recall Cls. F1 Cost (\$)
Ground Truth - 7.22 - - -
GPT-5 10 6.27 89.42 64.65 0.92
20 6.04 89.29 66.07 0.59
GPT-5-mini 10 5.54 92.17 68.80 0.16
20 5.19 90.66 67.52 0.10
Table 4: Comparison of claim-level extraction Density (claims per paragraph), recall, and classification F1 across models (low effort) and batch sizes.
@@
Grouped Retrieval F1 Cost (\$/1k)
87.25 12.84
10 90.91 3.65
10 83.10 0.95
20 91.61 3.46
20 87.25 0.95
Table 5: Ablation study on GPT-5-mini (low effort) comparing grouped verification and retrieval.
Table 5 presents an ablation study on grouped verification and retrieval using GPT-5-mini. Grouping multiple claims substantially reduces cost while maintaining strong verification accuracy. Without retrieval, increasing the group size to 20 achieves the highest F1 score (91.61) at a low cost. Retrieval further reduces the cost to \$0.95 per 1k claims but introduces a clear drop in accuracy, highlighting a trade-off between verification fidelity and efficiency. Overall, grouped verification without retrieval provides the best accuracy–cost balance, while retrieval-augmented settings are suitable for budget-constrained scenarios.
@@ 7 Conclusion
We propose DEER, a benchmark and evaluation framework for systematically assessing deep research reports. DEER builds a hierarchical taxonomy, instantiates it as 101 fixed rubrics, and provides task-specific expert guidance to improve the reliability of LLM-based judging. Beyond rubric scoring, DEER evaluates information use by tracing all claims back to their external information sources to verify evidence and quantifying evidence quality, enabling a more complete report-level assessment. Experiments show that deep research systems perform well on structure/style and information use, but remain limited in meeting expert-level requirements and producing analytically sound analyses. With its taxonomy, fixed rubrics, and quantitative metrics, DEER supports fine-grained diagnosis and systematic improvement beyond mere performance assessment.